Find all numbers with the property that when the number is added to itself the sum is the same as when the number is multiplied by itself. Could you help me with this Please? I can't figure out how to write the equation
Found 2 solutions by rapaljer, scott8148:
Answer by rapaljer(4667)

Let x = the number

x+x = x*x
2x= x^2

This is a quadratic equation, so you have to set it equal to zero. To keep the x^2 term positive, let's subtract 2x from each side, as follows:

Subtract 2x from each side:
0=x^2-2x

Factor out the common factor of x:
0=x(x-2)
x=0 or x-2=0
x=0 or x=2

If you need additional help solving "QUADRATIC EQUATIONS", please see my own website. Do a "Bing" or "Google" search for my last name "Rapalje". Look for "Rapalje Homepage" near the top of these searches. Then near the top of my Homepage, look for "Basic, Intermediate and College Algebra". Choose "Basic Algebra", "Chapter 2", "Section 2.09 Quadratic Equations". See also the "MATH IN LIVING COLOR" pages that go with this section.

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus


Answer by scott8148(6628)

n + n = n * n

2n = n^2

0 = n^2 - 2n

factoring ___ 0 = n (n - 2)

n = 0

n - 2 = 0 ___ n = 2

0 and 2 are the solutions